OpenStudy (anonymous):

Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (campbell_st):

if both points are on the line then find the gradient \[m = \frac{y_{2} - y_{1}}{x_{2} - x_{1}}\] which will give \[m = \frac{3 - 2}{1 - 0} ... m = 1\] the gradient of the line is m = 1 so the line is y = x + b use the point (1, 3) x = 1 and y = 3 and substitute into the equation above and allows you to find b 3 = 1 + b b = 2 then the equation is y = x + 2 in standard form the equation is x - y = -2